Please answer with proper calculations: no mean and sd required and not given.

value

8

10

6

11

1. The degrees of freedom for the test are:

2. The critical value of t for the test is:

3. The calculated t statistic is

4. Based on the statistical test you can

>reject the null hypothesis.

>fail to reject the null hypothesis.

>can not determine.

>No answer text provided.

5. The null hypothesis would be:

>there is no difference between the two population means.

>there is a difference between the two poulation means.

>there is only a positive idifference between the means.

>there is only a negative difference between the two populations.

6. The alternative hypothesis would be:

>test scores will be higher for the group using the new reading materials.

>test scores will be lower for the group using the new reading materials.

>there is no difference between the two population means.

>cannot determine.
Answer #1

Ansues Degree of freedom is df = N-1 at N=8 de = 8- 1 7 ot N de 10- 1 9 at N=6 de - 6-15 at N= 11 ds = 111 =10